The following day shows the relationship between the price and quantity demanded at four different prices for a product:Using the midpoints formula what is the price elasticity of demand between a) $11 and $9 ; (b)$9 and $7 ; (c) $7 and $5


1
Expert's answer
2022-03-23T16:22:25-0400

Since we don't have any data about quantity demanded at four different prices we can suggest they were following:

a) 16 and 24;

b) 24 and 32;

c) 32 and 40.

So we can calculate the price elasticity of demand:

a) "e_d=\frac{(24-16)/[(24+16)/2]}{(9-11)/[(9+11)/2]}=-2"

b)"e_d=\frac{(32-24)/[(32+24)/2]}{(7-9)/[(7+9)/2]}=-1.14"

c) "e_d=\frac{(40-32)/[(40+32)/2]}{(5-7)/[(5+7)/2]}=-0.667"